# Neutron star mountains supported by crustal lattice pressure

### v1

@@ 3 Thermal corrections to the electron gas pressure
The pressure in the outer part of a NS crust can be approximated as coming from a Fermi gas of relativistic electrons. The Fermi temperature $T_{\rm F}$ gives an estimate of the temperature at which thermal effects become important (Landau & Lifshitz (1969), Section 61):
[…]
@@
[…]
@@
In the limit of small $T/T_{\rm F}$, the total pressure can be divided into $P_{0}$, the zero-temperature piece, and an (analytically-calculable) thermal correction $P_{\rm therm}$. Using results from Landau & Lifshitz (1969), sections 61.4–61.6, we obtain
@@
Most relevant for mountains is the perturbation in pressure caused by a perturbation $\delta T$ in $T$:
$\frac{\delta P_{\rm therm}}{P_{0}}=4\pi^{2}\left(\frac{T}{T_{\rm F}}\right)^{2}\left(\frac{\delta T}{T}\right)=2.29\times 10^{-8}\,\left(\frac{\mu_{\rm e}}{3}\right)^{2/3}\frac{1}{\rho_{12}^{2/3}}T_{8}^{2}\left(\frac{\delta T/T}{1\%}\right),$ (12)
where we have parameterised the temperature perturbation as a $1\%$ change to the unperturbed value. We plot this fractional change in Figure 2, for an assumed uniform fractional temperature perturbation of $\delta T/T=1\%$. The fractional perturbation in pressure can be seen to vary from $\sim 10^{-5}$ in the outer crust, down to $\sim 10^{-10}$ near the crust-core transition. The implications of these results will be discussed in Section 5, where we shall compare the fractional perturbation in pressure due to each of the electrons, neutrons, and crustal lattice.
Figure 2: Fractional perturbation in pressure due to relativistic electrons for the Haensel & Zdunik (1990a, b) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state, as a function of density and temperature, for a temperature perturbation $\delta T/T=1\%$, computed using equation (12).
@@ 4 Thermal corrections to the neutron gas pressure
From Landau & Lifshitz (1969), section 57.3:
@@
In the case of small $T/T_{\rm F}$, the pressure can again be divided into zero-temperature and thermal pieces. Using results for sections 56–58 of Landau & Lifshitz (1969) we obtain
@@
Most relevant for mountains is the perturbation in pressure caused by a perturbation $\delta T$ in $T$:
$\frac{\delta P_{\rm therm}}{P_{0}}=\frac{5\pi^{2}}{4}\left(\frac{T}{T_{\rm F}}\right)^{2}\left(\frac{\delta T}{T}\right)=4.50\times 10^{-6}\frac{T_{8}^{2}}{(\rho_{12}X_{\rm n})^{4/3}}\left(\frac{\delta T/T}{1\%}\right).$ (18)
Like before, we plot this fractional change in Figure 3, again for an assumed uniform fractional temperature perturbation of $\delta T/T=1\%$. We use a lower density equal to that of neutron drip ($6.11\times 10^{11}$ g cm-3 for HZ90 and $4.38\times 10^{11}$ g cm-3 for BSk21), as this effect requires the existence of free neutrons. The magnitude of the fractional perturbation can be seen to range from $\sim 10^{-3}$ at the onset of neutron drip, down to $\sim 10^{-10}$ near the crust-core transition at low temperature.
Figure 3: Fractional perturbation in pressure due to non-relativistic neutrons for the Haensel & Zdunik (1990a, b) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state, as a function of density and temperature, for a temperature perturbation $\delta T/T=1\%$, computed using equation (18).
@@ 5 Thermal corrections to the lattice pressure
To decide which regime is applicable, we plot the plasma ion temperature for both the Haensel & Zdunik (1990a, b) EoS and the BSk21 Fantina et al. (2018, 2022) EoS in Figure 4. A typical NS in an LMXB has a temperature of a few times $10^{8}$ K, so that $T/T_{\rm pi}\sim 0.1$ over much of the crust, for both equations of state. This indicates that we are in the quantum regime, but not deeply so.
Figure 4: Plasma ion temperature $T_{\rm pi}$ for the Haensel & Zdunik (1990a, b) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state.
@@
For mountain building, it is more insightful to consider the perturbation in the thermal pressure, $\delta P_{\rm th}$, caused by a temperature perturbation $\delta T$, rather than $P_{\rm th}$ itself. From equations (28) and (30) we obtain
@@
$\frac{\delta P_{\rm th}}{P_{0}}=\frac{T}{P_{0}}\frac{dP_{\rm th}}{dT}\left(\frac{\delta T}{T}\right)$ (32)
to give the fractional change in pressure caused by a fractional temperature perturbation $\delta T/T$. Note that is it essentially this quantity that acts as the source term in the full perturbative analysis of the elastic equilibrium problem solved in Ushomirsky et al. (2000); see their equation (35) (but, of course, their perturbations were sourced by moving electron capture layers, not the thermal piece of the crustal EoS considered here.) A plot of this fractional change is given in Figure 5, for an assumed uniform fractional temperature perturbation of $\delta T/T=1\%$. The fractional perturbation in pressure can be seen to vary from $\sim 10^{-4}$ in the outer crust, down to $\sim 10^{-6}$ in the deep crust.
Figure 5: Fractional perturbation in pressure due to the crustal lattice for the Haensel & Zdunik (1990a, b) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state, as a function of density and temperature, as computed using the formalism described in Section 5, for a temperature perturbation $\delta T/T=1\%$.
@@
Before comparing the results we have obtained for pressure variations due to thermal lattice pressure with those of UCB for capture layers shifts (see Section 6), we will first make a quick comparison between the electron gas, neutron gas, and lattice pressure results given above. UCB found that much of their quadrupole was generated for densities of a few times $10^{12}$ g cm-3. A typical LMXB temperature is a few times $10^{8}$ K in this density range (see e.g. Hutchins & Jones 2023). If we look at this portion of the density-temperature parameter space of Figures 2, 3 and 5, we see that the fractional pressures perturbations for electrons, neutrons, and the crustal lattice, are $\sim 10^{-7}$, $\sim 10^{-6}$, and $\sim 10^{-6}$, respectively. It seems that the thermal component of crustal lattice pressure is competitive with the thermal fluctuations in the Fermi gas contributions, as well as having the advantage of being firmly tied in place by the crustal lattice itself.
@@ 6 Comparison between capture layer shifts and thermal crustal pressure
Having examined the effect of temperature on crustal lattice pressure, it is natural to try and compare its importance with the effective temperature dependence of the EoS that comes through the capture layer shifts considered in Ushomirsky et al. (2000). Rather than repeating their entire calculation for our neutron star model, we will make a rough comparison using their published results. UCB give plots showing results for the particular case of a star with a normal (i.e. not superfluid) core, accreting at a (rather high) rate of $0.5\dot{M}_{\rm Edd}$, half the Eddington accretion limit. Their temperature asymmetry is generated by an asymmetry in the local nuclear heating rate, but this is not of importance for the purposes of this comparison, only the temperature perturbation itself is needed to compute the corresponding pressure perturbation. This temperature perturbation is given in the upper left panel of their Figure 8, and can be seen to be at the few percent level, specifically $\delta T/T\sim 3\%$ over the density range where their capture layers occur.
@@
The corresponding fractional perturbations in pressure, as defined by our equation (32) for this $\sim 3\%$ temperature perturbation, are then given in the lower panels of UCB’s Figures 10-12. In their Figure 10 they give results for a shallow capture layer with capture energy $E_{\rm cap}=23$ MeV. In their Figure 11 they give results for a deeper capture layer with $E_{\rm cap}=42$ MeV, corresponding to one of the deepest capture layers in the computation of Haensel & Zdunik (1990b). In their Figure 12 they gives results for a capture layer with $E_{\rm cap}=95$ MeV. As we noted in Section 2, no such capture layer was found in the computations of Haensel & Zdunik (1990b) or Fantina et al. (2018). Given this, we will focus our comparison on the second of these capture layers.
@@
For this $E_{\rm cap}=42$ MeV capture layer, Table 2 of Haensel & Zdunik (1990b) shows that $\rho\approx 9.0\times 10^{12}$ g cm-3 and $P\approx 1.2\times 10^{31}$ erg cm-3. For the particular accreting star that UCB consider, their Figure 7 indicates that $T\approx 6\times 10^{8}$ K. Note that, from the left panel of our Figure 5, we have $\delta P_{\rm th}/P_{0}\approx 3\times 10^{-6}$ for this combination of density and temperature and temperature perturbation $\delta T/T\approx 3\%$, assuming (for consistency) the EoS of Haensel & Zdunik (1990a, b).
In contrast, UCB, from the bottom panel of their Figure 11, peaks at around $\delta P_{\rm th}/P\sim 10^{-2}$. This indicates that within their capture layer, the thermal component of the pressure is very important: a percent-level temperature variation is resulting in a percent-level pressure perturbation. However, the function is extremely sharply peaked in radius, with a full-width half maximum $\sim 10^{2}$ cm. Averaging over the crustal thickness of $\sim 10^{5}$ cm, the radially averaged fractional pressure perturbation is $\langle\delta P_{\rm th}/P\rangle\sim 10^{-5}$. Allowing for the existence of $\sim 10$ such capture layers gives $\langle\delta P_{\rm th}/P\rangle\sim 10^{-4}$.
@@ 7 Summary and Discussion
Gravitational waves from thermoelastic mountains may play a role in determining the spin frequencies of neutron stars in LMXBs (Bildsten, 1998). The leading candidate for producing such deformations relies on the existence of electron capture layers with high threshold energies, deep in the crust (Ushomirsky et al., 2000). As we described in Section 2, modern EoS calculations do not provide support for the existence of such deep layers (Fantina et al., 2018, 2022), with a recent study even casting doubt over the existence of there being any capture layers at densities above neutron drip (Gusakov & Chugunov, 2020; Potekhin et al., 2023).
@@
We therefore advance such crustal lattice pressure asymmetries as worthy of further investigation. We are carrying out such a study ourselves, with the star’s magnetic field providing the source of the temperature asymmetry, building on the work presented in Osborne & Jones (2020) and Hutchins & Jones (2023). We will present this in a separate study.

### v2

@@ 3 Thermal corrections to the electron gas pressure
The pressure in the outer part of a NS crust can be approximated as coming from a Fermi gas of relativistic electrons. The Fermi temperature $T_{\rm F}$ gives an estimate of the temperature at which thermal effects become important (Landau and Lifshitz (1969), Section 61):
$T_{\rm F}=(3\pi^{2})^{1/3}\frac{\hbar c}{k_{\rm B}}\left(\frac{\rho}{\mu_{\rm e}m_{\rm u}}\right)^{1/3}=4.15\,\times\,10^{11}\,{\rm K\,}\rho_{12}^{1/3}\left(\frac{3}{\mu_{\rm e}}\right)^{1/3}.$ (8)
@@
$\frac{T}{T_{\rm F}}=T\frac{1}{(3\pi^{2})^{1/3}}\frac{k_{\rm B}}{\hbar c}\left(\frac{\mu_{\rm e}m_{\rm u}}{\rho}\right)^{1/3}=2.42\,\times\,10^{-4}\,T_{8}\left(\frac{\mu_{\rm e}}{3}\right)^{1/3}\frac{1}{\rho_{12}^{1/3}}.$ (9)
@@
In the limit of small $T/T_{\rm F}$, the total pressure can be divided into $P_{0}$, the zero-temperature piece, and an (analytically-calculable) thermal correction $P_{\rm therm}$. Using results from Landau and Lifshitz (1969), sections 61.4–61.6, we obtain
@@
Most relevant for mountains is the perturbation in pressure caused by a perturbation $\delta T_{22}$ in $T$:
$\frac{\delta P_{\rm therm}}{P_{0}}=4\pi^{2}\left(\frac{T}{T_{\rm F}}\right)^{2}\left(\frac{\delta T_{22}}{T}\right)=2.29\times 10^{-8}\,\left(\frac{\mu_{\rm e}}{3}\right)^{2/3}\frac{1}{\rho_{12}^{2/3}}T_{8}^{2}\left(\frac{\delta T_{22}/T}{1\%}\right),$ (12)
where we have parameterised the temperature perturbation as a $1\%$ change to the unperturbed value. We plot this fractional change in Figure 2, for an assumed uniform fractional temperature perturbation of $\delta T_{22}/T=1\%$. The fractional perturbation in pressure can be seen to vary from $\sim 10^{-5}$ in the outer crust, down to $\sim 10^{-10}$ near the crust-core transition. The implications of these results will be discussed in Section 5, where we shall compare the fractional perturbation in pressure due to each of the electrons, neutrons, and crustal lattice.
Figure 2: Fractional perturbation in pressure due to relativistic electrons for the Haensel and Zdunik (1990b, a) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state, as a function of density and temperature, for a temperature perturbation $\delta T_{22}/T=1\%$, computed using equation (12).
@@ 4 Thermal corrections to the neutron gas pressure
From Landau and Lifshitz (1969), section 57.3:
@@
In the case of small $T/T_{\rm F}$, the pressure can again be divided into zero-temperature and thermal pieces. Using results for sections 56–58 of Landau and Lifshitz (1969) we obtain
@@
Most relevant for mountains is the perturbation in pressure caused by a perturbation $\delta T_{22}$ in $T$:
$\frac{\delta P_{\rm therm}}{P_{0}}=\frac{5\pi^{2}}{4}\left(\frac{T}{T_{\rm F}}\right)^{2}\left(\frac{\delta T_{22}}{T}\right)=4.50\times 10^{-6}\frac{T_{8}^{2}}{(\rho_{12}X_{\rm n})^{4/3}}\left(\frac{\delta T_{22}/T}{1\%}\right).$ (18)
Like before, we plot this fractional change in Figure 3, again for an assumed uniform fractional temperature perturbation of $\delta T_{22}/T=1\%$. We use a lower density equal to that of neutron drip ($6.11\times 10^{11}$ g cm-3 for HZ90 and $4.38\times 10^{11}$ g cm-3 for BSk21), as this effect requires the existence of free neutrons. The magnitude of the fractional perturbation can be seen to range from $\sim 10^{-3}$ at the onset of neutron drip, down to $\sim 10^{-10}$ near the crust-core transition at low temperature.
Figure 3: Fractional perturbation in pressure due to non-relativistic neutrons for the Haensel and Zdunik (1990b, a) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state, as a function of density and temperature, for a temperature perturbation $\delta T_{22}/T=1\%$, computed using equation (18).
@@ 5 Thermal corrections to the lattice pressure
To decide which regime is applicable, we plot the plasma ion temperature for both the Haensel and Zdunik (1990b, a) EoS and the BSk21 Fantina et al. (2018, 2022) EoS in Figure 4. A typical NS in an LMXB has a temperature of a few times $10^{8}$ K, so that $T/T_{\rm pi}\sim 0.1$ over much of the crust, for both equations of state. This indicates that we are in the quantum regime, but not deeply so.
Figure 4: Plasma ion temperature $T_{\rm pi}$ for the Haensel and Zdunik (1990b, a) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state.
@@
For mountain building, it is more insightful to consider the perturbation in the thermal pressure, $\delta P_{\rm th}$, caused by a temperature perturbation $\delta T_{22}$, rather than $P_{\rm th}$ itself. From equations (28) and (30) we obtain
@@
$\frac{\delta P_{\rm th}}{P_{0}}=\frac{T}{P_{0}}\frac{dP_{\rm th}}{dT}\left(\frac{\delta T_{22}}{T}\right)$ (32)
to give the fractional change in pressure caused by a fractional temperature perturbation $\delta T_{22}/T$. Note that is it essentially this quantity that acts as the source term in the full perturbative analysis of the elastic equilibrium problem solved in Ushomirsky et al. (2000); see their equation (35) (but, of course, their perturbations were sourced by moving electron capture layers, not the thermal piece of the crustal EoS considered here.) A plot of this fractional change is given in Figure 5, for an assumed uniform fractional temperature perturbation of $\delta T_{22}/T=1\%$. The fractional perturbation in pressure can be seen to vary from $\sim 10^{-4}$ in the outer crust, down to $\sim 10^{-6}$ in the deep crust.
Figure 5: Fractional perturbation in pressure due to the crustal lattice for the Haensel and Zdunik (1990b, a) (left) and BSk21 (Fantina et al. 2018, 2022; right) equations of state, as a function of density and temperature, as computed using the formalism described in Section 5, for a temperature perturbation $\delta T_{22}/T=1\%$.
@@
Before comparing the results we have obtained for pressure variations due to thermal lattice pressure with those of UCB for capture layers shifts (see Section 6), we will first make a quick comparison between the electron gas, neutron gas, and lattice pressure results given above. UCB found that much of their quadrupole was generated for densities of a few times $10^{12}$ g cm-3. A typical LMXB temperature is a few times $10^{8}$ K in this density range (see e.g. Hutchins and Jones 2023). If we look at this portion of the density-temperature parameter space of Figures 2, 3 and 5, we see that the fractional pressures perturbations for electrons, neutrons, and the crustal lattice, are $\sim 10^{-7}$, $\sim 10^{-6}$, and $\sim 10^{-6}$, respectively. It seems that the thermal component of crustal lattice pressure is competitive with the thermal fluctuations in the Fermi gas contributions, as well as having the advantage of being firmly tied in place by the crustal lattice itself.
@@ 6 Comparison between capture layer shifts and thermal crustal pressure
Having examined the effect of temperature on crustal lattice pressure, it is natural to try and compare its importance with the effective temperature dependence of the EoS that comes through the capture layer shifts considered in Ushomirsky et al. (2000). Rather than repeating their entire calculation for our neutron star model, we will make a rough comparison using their published results.
Specifically, we will, for a given temperature perturbation $\delta T_{22}$, compare the sizes of the corresponding pressure perturbations $\delta P$, for our crustal lattice pressure perturbations and UCB’s shifting capture layers. The justification for using this simple method of comparison lies in the fact that the relevant set of ODEs (equations 43(a)–43(d) of UCB) are linear in both the crust’s elastic displacement and in the source term itself, implying that larger source terms automatically produce larger elastic responses. Of course, the exact radial profile of the perturbation would differ in the two cases, the effect of which can only be quantified with a full numerical solution (which we will present in a subsequent work).
In particular, the source terms in UCB are very sharply peaked at the capture layers that divide shells of different atomic number, whereas our source terms as smooth, being directly obtained from the star’s thermal profile, as per our equation (32). In making our rough comparisons below, we are implicitly assuming that this sharpness does not introduce any qualitatively different features in the elastic response of the crust and the resultant density perturbations, or any sort of “screening” effect where perturbations at different densities nearly cancel one another. Such a screening effect has been seen in the tidal responses of two layer quark stars (Lau et al., 2019), but there is no reason to expect it to apply in our case.
UCB give plots showing results for the particular case of a star with a normal (i.e. not superfluid) core, accreting at a (rather high) rate of $0.5\dot{M}_{\rm Edd}$, half the Eddington accretion limit. Their temperature asymmetry is generated by an asymmetry in the local nuclear heating rate, but this is not of importance for the purposes of this comparison, only the temperature perturbation itself is needed to compute the corresponding pressure perturbation. This temperature perturbation is given in the upper left panel of their Figure 8, and can be seen to be at the few percent level, specifically $\delta T_{22}/T\sim 3\%$ over the density range where their capture layers occur.
@@
The corresponding fractional perturbations in pressure, as defined by our equation (32) for this $\sim 3\%$ temperature perturbation, are then given in the lower panels of UCB’s Figures 10-12. In their Figure 10 they give results for a shallow capture layer with capture energy $E_{\rm cap}=23$ MeV. In their Figure 11 they give results for a deeper capture layer with $E_{\rm cap}=42$ MeV, corresponding to one of the deepest capture layers in the computation of Haensel and Zdunik (1990a). In their Figure 12 they gives results for a capture layer with $E_{\rm cap}=95$ MeV. As we noted in Section 2, no such capture layer was found in the computations of Haensel and Zdunik (1990a) or Fantina et al. (2018). Given this, we will focus our comparison on the second of these capture layers.
@@
For this $E_{\rm cap}=42$ MeV capture layer, Table 2 of Haensel and Zdunik (1990a) shows that $\rho\approx 9.0\times 10^{12}$ g cm-3 and $P\approx 1.2\times 10^{31}$ erg cm-3. For the particular accreting star that UCB consider, their Figure 7 indicates that $T\approx 6\times 10^{8}$ K. Note that, from the left panel of our Figure 5, we have $\delta P_{\rm th}/P_{0}\approx 3\times 10^{-6}$ for this combination of density and temperature and temperature perturbation $\delta T_{22}/T\approx 3\%$, assuming (for consistency) the EoS of Haensel and Zdunik (1990b, a). In contrast, the pressure perturbation of UCB, from the bottom panel of their Figure 11, peaks at around $\delta P_{\rm th}/P\sim 10^{-2}$. This indicates that within their capture layer, the thermal component of the pressure is very important: a percent-level temperature variation is resulting in a percent-level pressure perturbation.
However, the function is extremely sharply peaked in radius, with a full-width half maximum $\sim 10^{2}$ cm. This sharpness stands in contrast to the smooth variation in pressure with radius that holds for the crustal lattice pressure perturbations; observe the smooth variation of $\delta P$ with density in Figure 5. Physically, this is related to the fact that for UCB, the pressure perturbation is coming from the change in the (spatially thin) capture layer, whereas for us it comes from the crustal lattice pressure, which varies smoothly over the crust.
Given this, in comparing UCB’s pressure perturbation with ours, we first carry out an average with respect to radius, the radius being the relevant independent variable on the set of linear ODEs that determine the elastic response of the star (again, see equations 43(a)–43(d) of UCB). Averaging over the crustal thickness of $\sim 10^{5}$ cm, the radially averaged fractional pressure perturbation is $\langle\delta P_{\rm th}/P\rangle\sim 10^{-5}$. Allowing for the existence of $\sim 10$ such capture layers gives $\langle\delta P_{\rm th}/P\rangle\sim 10^{-4}$.
@@
As a caveat, we remind the reader that in a real neutron star, the pure $l=m=2$ deformations considered here will exist alongside other multipole contributions, dependent upon the history of the crust since formation. The entire history of strain, fracture and plastic flow will be relevant. These extra contributions to the strain will add, and it is the sum of these that will ultimately be limited by the finite strength of the crust. See the material surrounding equation (34) of UCB for more discussion of this.
@@ 7 Summary and Discussion
Gravitational waves from thermoelastic mountains may play a role in determining the spin frequencies of neutron stars in LMXBs (Bildsten, 1998). The leading candidate for producing such deformations relies on the existence of electron capture layers with high threshold energies, deep in the crust (Ushomirsky et al., 2000). As we described in Section 2, modern EoS calculations do not provide support for the existence of such deep layers (Fantina et al., 2018, 2022), with a recent study even casting doubt over the existence of there being any capture layers at densities above neutron drip (Gusakov and Chugunov, 2020; Potekhin et al., 2023).
@@
We therefore advance such crustal lattice pressure asymmetries as worthy of further investigation. We are carrying out such a study ourselves, with the star’s magnetic field providing the source of the temperature asymmetry, building on the work presented in Osborne and Jones (2020) and Hutchins and Jones (2023). We will present this in a separate study.